34. Sailright Inc. manufactures and sells sailboards. Management believes that the price elasticity of demand is – 3.0. Currently, boards are priced at $500 and the quantity demanded is 10,000 per year.
(a) If the price is increased to $600, how many sailboards will the company be able to sell each year?
(b) How much will total revenue change as a result of the price increase?
1
Expert's answer
2020-11-19T06:45:55-0500

(a) If the price is increased to $600, how many sailboards will the company be able to sell each year?

Change in price (600-500)/500 = "\frac{100}{500} = 0.2 (20\%)"

Elasticity of demand = -3

Change in quantity demanded = "20\times-3 = -60"

Quantity demanded = "10,000\times60\% = 6000"

(b) How much will total revenue change as a result of the price increase?

Total revenue = Price * Quantity

TR = "600\times6,000" = $ 3,600,000